Q. Choose the correct answer.
Q. If R1, R2 and R3 are radii of circles x2+y2-2x+2y-2=0, x2+y2+2x-2y-7=0 and x2+y2+4x-2y-11=0 respectively, then
a. R1>R2>R3
b. R2>R1>R3
c. R3>R2>R1
d. R3>R1>R2
1
Expert's answer
2019-07-15T09:09:13-0400

General equation of circle is

x2+y2+2gx+2fy+c=0

And radius is given by ( g2+f2-c)1/2

R1=(12+12+2)1/2=2

R2=(12+12+7)1/2=3

R3=(22+12+11)1/2=4

Hence R3 >R2>R1

Option C is correct.